# ReWiND: Language-Guided Rewards Teach Robot Policies without New Demonstrations

### v1

@@ 1 Introduction
Our core contribution is in designing ReWiND’s reward model to capture three key properties outlined earlier: dense feedback, generalization, and robustness. First, to provide dense, informative feedback, we design a cross-modal sequential aggregator that predicts progress within demonstration videos. Progress prediction offers a densely supervised training signal that naturally translates into rewards. We also introduce video rewinding to automatically generate failure trajectories from successful demos, allowing ReWiND to provide dense feedback even when the policy is making mistakes. Then, to encourage generalization across unseen tasks and robustness to diverse behaviors, we train the cross-modal sequential aggregatorwith pre-trained vision and language embeddings, selectively applied positional embeddings, and diverse robotics data from Open-X [16]. Focusing on these properties enables ReWiND rewards to extrapolate to novel visual and language inputs.
@@ 2 Related Works
Other approaches use pre-trained vision models to derive rewards from visual observations [35, 36, 37, 38, 39, 6, 14, 13, 40, 41, 42, 43, 44, 45]. Among these, RoboCLIP [6], LIV [10], VLC [13], and GVL [11]—like ReWiND—reward unseen robot manipulation tasks directly from language without additional target-task demos or online tuning. We show in Section 4.1 that these baselines underperform ReWiND in rewarding policies in our limited-data setting. Most similar to ReWiND, Foundation Actor-Critic (FAC) [46] enables efficient RL from language via potential-based shaping rewards from a pre-trained VLM. However, FAC depends on predefined policy priors (e.g., code-based primitives from LLMs), whereas ReWiND learns them through offline RL on non-target tasks.
@@ 3.1 Learning a Reward Function
Our primary objective for reward prediction is regressing directly to per-frame progress within an observation sequence conditioned on instruction . Unlike prior methods using relative targets [41, 13], our progress-based objective provides fixed targets that are more stable to train on, and translates directly into a dense, -normalized reward for policy training. To ensure robustness against mismatched observations and instructions, we also sample unrelated observation sequences and train to predict zero progress. Our reward prediction loss is:
@@ 3.2 Policy Learning
Learning Online. To learn a new task online, ReWiND only requires a language description of the task, . ReWiND rolls out and fine-tunes it on rewards coming from . Like prior work [13, 41], we assume access to a success signal during online RL. We use this signal to give bonuses similar to in pre-training.222Success bonuses can come from a human supervisor [55], learned function [23], or LLM [46]. Our experiments assume a human supervisor because manual resets are required regardless. While we could threshold outputs to automatically determine success, unseen evaluation task reward ranges can vary, rendering this approach ineffective. Future work could integrate ReWiND with methods reducing human resets [42, 56] and automatic success detectors for truly autonomous RL. Our online rewards are:
@@ 4.1 Q1: What Makes a Good Reward Function?
Next, we evaluate how consistently rewards reflect progress over time in successful, unseen demonstrations. We report Pearson correlation () of each model’s reward against time, and Spearman’s rank correlation (), which, unlike , captures monotonicity regardless of linearity. As shown in Table 1(a), ReWiND again outperforms all baselines—achieving a 30% relative improvement in and 27% in over the best alternative (VLC).
@@
We evaluate each dataset’s relative alignment ranking (measured by Spearman’s ) with each reward model. For example, for a given task, if the average reward for a failure video is 0.1, a near-success video is 0.5, and success video is 0.9, then the rankings would be 1, 2, 3, respectively, where 3 corresponds to the best ranking. Thus, over the rankings tells us how often the videos are correctly ranked. We report the ranking in Table 1(b). We also report the average difference between rewards for success with near-success and near-success with failure videos. Overall, likely due to video rewinding, ReWiND has a relative 74% improvement in reward order and 58% improvement in reward differences over the best baseline, LIV-FT. Additionally, we qualitatively demonstrate how these rankings translate into policy rollout rewards in Appendix Figure 8 by plotting per-frame reward curve predictions of ReWiND against reward baselines for an unsuccessful policy rollout.
@@ 4.2 Q2: Learning New Tasks with RL
Simulation. We use the Meta-World simulation benchmark [62], where we pre-train reward models and policies on 20 tasks, each with 5 per-task demos collected from a scripted policy. We evaluate on 8 unseen tasks in Meta-World, chosen for reasonable initial policy rollout behaviors, across 3 seeds each. We compare ReWiND against the 2 language-conditioned reward model baselines that performed best in reward alignment (VLC) and policy rollout rankings (LIV-FT) from the reward analysis in Section 4.1. We also compare against Sparse, which pre-trains and fine-tunes on only the sparse success reward bonus, and Pre-train, which pre-trains on sparse reward and is evaluated zero-shot on new tasks. All baselines are image, proprioception (), and language conditioned. Each method uses the same policy pre-training and RL procedure as ReWiND as outlined in Section 3.2, and is trained online for 100k timesteps. See Appendix B for environment and policy training details.
As recommended by Agarwal et al. [64], we report the interquartile mean (IQM) and 95% confidence intervals computed over all task success rates at 100k environment steps in Figure 5. Sparse reward fine-tuning and Pre-train (no fine-tuning) result in near-zero success rates, highlighting the difficulty of image-based new task learning under limited data. In fact, Sparse reward fine-tuning, which relies purely on a sparse success bonus, performs worse than Pre-train after fine-tuning. Meanwhile, ReWiND achieves an IQM success rate of 79%, a 97.5% improvement over the best baseline, VLC, demonstrating that ReWiND effectively enables the policy to learn new tasks in Meta-World. These results are well-aligned with our reward analysis in Section 4.1, demonstrating how they correlate with policy learning performance. ReWiND is also more sample-efficient at timesteps less than 100k; see extended discussion in Section D.2 and sample efficiency curves in Figure 14(i).
@@
We conduct real-world tabletop manipulation experiments with a bimanual Koch v1.1 robot arm setup [65]. We use 5 demos to train the reward function, but 10 for the policy, as we found policy learning to be a bottleneck on this difficult robot embodiment. Across five tasks, we demonstrate in Figure 6 that an hour of real-world RL with ReWiND improves the success rate over the base pre-trained policy from an average 12% success rate to 68%, a 5 improvement. Meanwhile, VLC only improves from 8% to 10%—ReWiND outperforms VLC, the best simulation baseline, by 6.7. RL for an hour of real-world experiment time corresponds to 50k environment steps with our parallelized codebase that trains the policy while an older checkpoint gathers data in the environment to avoid any training wait time. We select diverse tasks that demonstrate real-world improvement based on generalization metrics defined in prior work [47, 48] on: an in-distribution task, separate the blue and orange cups; an in-distribution difficult task, fold the blue towel; an unseen task in terms of large amounts of visual clutter, open the red trash bin; an unseen task in terms of spatial relationships between objects requiring new action sequences, put the orange cup on the red plate; and an unseen task in terms of language input, put the fruit-colored object in the box. Overall, ReWiND enables real-world reinforcement learning on unseen tasks without requiring new demonstrations, improving over the pretrained pre-trained policy, and outperforms the best baseline from simulation, VLC. See Section C.1 for real-world experiment details and Figure 12 for policy rollout examples.
@@ 4.3 Q3: Ablation Study
Augmentation. Eliminating video rewinding (Video Rewind) degrades rollout ranking performance (Table 2b), showing that rewinding helps distinguish failed rollouts as intended. This variant performs similarly to the single-image LIV-FT baseline in Table 1, indicating that video rewinding more effectively captures the temporal information in the videos. Similarly, removing instruction generation (Instruction Generation) reduces performance on language input robustness (Table 2c), confirming that LLM-generated instructions enhance robustness to diverse inputs.
Architecture. Adding full positional embeddings (Full Pos. Embeds) improves unseen demo alignment (Table 2a) but worsens rollout ranking (Table 2b), likely due to overfitting—where the model learns to predict increasing rewards regardless of input. To avoid overfitting, the main ReWiND model uses only first-frame positional embeddings (Section 3.1.2).
@@ 5 Limitations
However, the best way to fine-tune these models with online rewards remains an open challenge [68, 69]. One bottleneck is simply that, even with low-rank adaptation techniques that prior work found to help train large policy architectures more efficiently [70, 69, 66], fine-tuning these models takes a lot of compute and real-world time that makes real-world online learning with reward difficult. We plan to investigate blending ReWiND approaches with such policy architectures in the future.
@@
Another limitation comes from requiring success detection for the reward bonus and terminating policy rollouts upon success. We add a success bonus (detailed in Section 3.2) to account for potential noisy rewards and imperfect success detection by the reward model, given that a human is already monitoring to reset the environment, and we subsequently terminate the rollout upon success. Methods such as those introduced by Ye et al. [46], Zhou et al. [73], Yang et al. [42], which utilize VLMs as success detectors, can remove the need for human supervision during the online phase of ReWiND when combined with reset-free RL. In future work, we plan to investigate the combination of ReWiND with reset-free approaches and automatic success detection for truly autonomous learning.
@@ Acknowledgments
This work was supported in part by a grant from the Army Research Lab (ARL) Army AI Innovations Institute (A2I2), award number W911NF-23-2-0010. The claims and findings of this work do not necessarily represent the views of the ARL. This work was also supported in part by a grant from the DARPA Friction and Accountability in Conversational Transactions (FACT) program.
@@ A.1.1 Open-X Dataset
We select a subset of datasets from the Open-X Dataset [16]. The subset includes Bridge-V2 [74], BC-Z [75], Fractal [1], CLVR Jaco Play [76], Berkeley Autolab UR5 [77], Berkeley Fanuc Manipulation [78], CMU Stretch [79, 80], Stanford Hydra [81], UCSD Kitchen [82], Austin BUDS [83], and Austin Sirius [84]. These datasets were selected for their high-quality, task-oriented manipulation trajectories (i.e., no play data or extremely high-level annotations). These datasets provide around 350k trajectories and 58k total unique task annotations. To ensure meaningful trajectories for training the ReWiND reward model, we postprocess the data to remove trajectories with less than 5 timesteps. We subsample the videos in the datasets to 16 frames for reward model training, as we did not see a noticeable benefit from training it with longer videos.
@@ A.1.2 Reward Function
The cross-modal sequential aggregator takes these tokens as input and produces a per-image embedding used by an MLP to produce per-timestep rewards. The cross-modal sequential aggregator is a causally masked transformer (PyTorch nn.TransformerEncoder) composed of 4 layers, each with 8 heads with a combined hidden dimension of 2048. We add a learnable positional embedding to only the first frame of the video sequence embedding. In the ReWind reward function training phase, we trained 2k steps for Meta-World and 10k steps for Real-World robot experiments, with a batch size of 1024. Each batch includes 80% data from and 20% target environment data from . Each video in the batch has an 80% probability of having video rewind augmentation, and independently, a 20% percent probability of having a mismatched video-language pairing with 0 progress target (see Section 3.1). In order to better policy execution videos that look close to success, 10% of the rewound videos will only have their last 3 frames rewound. No extensive tuning was performed on these per-sample rewind and mismatch probabilities; they were heuristically chosen during initial small-scale experimentation and then fixed for all experiments.

### v2

@@ 1 Introduction
Our core contribution is in designing ReWiND’s reward model to capture three key properties outlined earlier: dense feedback, generalization, and robustness. First, to provide dense, informative feedback, we design a cross-modal sequential aggregator that predicts progress within demonstration videos. Progress prediction offers a densely supervised training signal that naturally translates into rewards. Importantly, we also introduce video rewind to automatically generate failure trajectories from successful ones, allowing ReWiND to provide dense feedback even when the policy is making mistakes. Then, to encourage generalization to unseen tasks and robustness to diverse behaviors, we train the sequential aggregator with pre-trained vision and language encoders, selectively applied positional embeddings, and diverse robotics data from Open-X [16]. Focusing on the above three properties enables ReWiND rewards to extrapolate to novel visual and language inputs.
@@ 2 Related Works
Other approaches use pre-trained vision models to derive rewards from visual observations [35, 36, 37, 38, 39, 6, 14, 13, 40, 41, 42, 43, 44, 45, 46]. Among these, RoboCLIP [6], LIV [10], VLC [13], VICTOR [46], and GVL [11]—like ReWiND—reward unseen manipulation tasks directly from language without sub-task annotations or target-task demos. We show in Section 4.1 that these baselines underperform ReWiND in rewarding policies in our limited-data setting. Similar to ReWiND, Foundation Actor-Critic [47] enables efficient RL from language via potential-based shaping rewards from a pre-trained VLM. However, it depends on predefined policy priors (e.g., code-based primitives from LLMs), whereas ReWiND learns them through offline RL on non-target tasks.
@@ 3.1 Learning a Reward Function
Our primary objective for reward prediction is regressing directly to per-frame progress within an observation sequence conditioned on instruction . Unlike prior methods using relative targets [41, 13], our progress-based objective provides stable, fixed targets, and translates directly into a dense, -normalized reward for policy training. To ensure robustness against mismatched observations and instructions, we also sample unrelated observation sequences and train to predict zero progress. Our reward prediction loss is:
@@ 3.2 Policy Learning
Learning Online. To learn a new task online, ReWiND only requires a language description of the task, . ReWiND rolls out and fine-tunes it on rewards coming from . Like prior work [13, 41], we assume access to a success signal during online RL. We use this signal to give bonuses similar to pre-training.222Success bonuses can come from a human supervisor [56], learned function [23], or LLM [47]. Our experiments assume a human supervisor because manual resets are required regardless. While we could threshold outputs to automatically determine success, unseen evaluation task reward ranges can vary, rendering this approach ineffective. Future work could integrate ReWiND with methods reducing human resets [42, 57] and automatic success detectors for truly autonomous RL. Our online rewards are:
@@ 4.1 Q1: What Makes a Good Reward Function?
Next, we evaluate how consistently rewards reflect progress over time in successful, unseen demonstrations. We report Pearson correlation () of each model’s per-frame reward against time and also Spearman’s rank correlation (), which, unlike , captures monotonicity regardless of linearity. As shown in Table 1(a), ReWiND again outperforms all baselines—achieving a 30% relative improvement in and 27% in over the best alternative (VLC).
@@
We evaluate each dataset’s relative alignment ranking (measured by Spearman’s ) with each reward model. For example, for a given task, if the average reward for a failure video is 0.1, a near-success video is 0.5, and success video is 0.9, then the rankings would be 1, 2, 3, respectively, where 3 corresponds to the best ranking. Thus, over rankings tells us how often the videos are correctly ranked. We report the ranking in Table 1(b). We also report the average difference between rewards for success with near-success and near-success with failure videos. Overall, due to video rewind, ReWiND has a relative 74% improvement in reward order and 58% improvement in reward difference over the best baseline, LIV-FT. Additionally, we qualitatively demonstrate how these rankings translate into policy rollout rewards in Appendix Figure 9 by plotting reward predictions of ReWiND against reward baselines for an unsuccessful policy rollout.
@@ 4.2 Q2: Learning New Tasks with RL
Simulation. We use the MetaWorld simulation benchmark [63], where we pre-train reward models and policies on 20 tasks, each with 5 per-task demos collected from a scripted policy. We evaluate on 8 unseen tasks in MetaWorld, chosen for reasonable initial policy rollout behaviors, across 3 seeds each. We compare ReWiND against the 2 language-conditioned reward model baselines that performed best in reward alignment (VLC) and policy rollout rankings (LIV-FT) from the reward analysis in Section 4.1. We also compare against Sparse, which pre-trains and fine-tunes on only the sparse success reward bonus, and Pre-train, which pre-trains on sparse reward and is evaluated zero-shot on new tasks. All baselines are image, proprio (), and language conditioned. Each method uses the same pre-training and RL procedure as ReWiND (Section 3.2), and is trained online for 100k timesteps. See Appendix B for more experiment details.
As recommended by Agarwal et al. [65], we report the interquartile mean (IQM) and 95% confidence intervals computed over all task success rates at 100k environment steps in Figure 5. Sparse reward fine-tuning and Pre-train (no fine-tuning) result in near-zero success rates, highlighting the difficulty of image-based new task learning under limited data. In fact, Sparse reward fine-tuning, which relies purely on a sparse success bonus, performs worse than Pre-train after fine-tuning. Meanwhile, ReWiND achieves an IQM success rate of 79%, a 97.5% improvement over the best baseline, VLC, demonstrating that ReWiND effectively enables the policy to learn new tasks in MetaWorld. These results are well-aligned with our reward analysis in Section 4.1, demonstrating how they correlate with policy learning performance. ReWiND is also more sample-efficient at timesteps less than 100k; see extended discussion in Section D.2 and sample efficiency curves in Figure 11.
@@
We conduct real-world tabletop manipulation experiments with a bimanual Koch v1.1 robot arm setup [66]. We use 5 demos to train the reward function, but 10 for the policy, as we found policy learning to be a bottleneck on this difficult robot embodiment. Across five tasks, we demonstrate in Figure 6 that an hour of real-world RL with ReWiND improves the success rate over the base pre-trained policy from an average 12% success rate to 68%, a 5 improvement. Meanwhile, VLC only improves from 8% to 10%—ReWiND outperforms VLC, the best simulation baseline, by 6.7. RL for an hour of real-world experiment time corresponds to 50k environment steps with our parallelized codebase that trains the policy while an older checkpoint gathers data in the environment to avoid any training wait time. We select diverse tasks that demonstrate real-world improvement based on generalization metrics defined in prior work [48, 49] on: an in-distribution task, separate the blue and orange cups; an in-distribution difficult task, fold the blue towel; an unseen task in terms of large amounts of visual clutter, open the red trash bin; an unseen task in terms of spatial relationships between objects requiring new action sequences, put the orange cup on the red plate; and an unseen task in terms of language input, put the fruit-colored object in the box. Overall, ReWiND enables real-world reinforcement learning on unseen tasks without requiring new demonstrations, improving over the pre-trained policy, and outperforms the best baseline from simulation, VLC. See Section C.1 for real-world experiment details and Figure 13 for policy rollout examples.
@@
| ReWiND Variant | Rank Order Corr. | Avg. Robustness Corr. | Policy Succ. Rate |
| 5 Task | 0.76 | 0.86 | 0.52 |
| 10 Task | 0.90 | 0.88 | 0.47 |
| + Full PE | 0.71 | 1.05 | 0.69 |
| - Video Rewind | 0.87 | 0.89 | 0.67 |
| ReWiND | 1.00 | 1.00 | 1.00 |
@@ 4.3 Q3: Ablation Study
Augmentation. Eliminating video rewinding (Video Rewind) degrades rollout ranking performance (Table 2b), showing that rewinding helps distinguish failed rollouts as intended. This variant performs similarly to the single-image LIV-FT baseline in Table 1, indicating that video rewinding more effectively captures the temporal information in the videos. It also reduces policy learning performance by 33% as seen in Table 3. Similarly, removing instruction generation (Instruction Generation) reduces performance on language input robustness (Table 2c), confirming that LLM-generated instructions enhance robustness to diverse inputs.
Architecture. Adding full positional embeddings (Full Pos. Embeds) improves unseen demo alignment (Table 2a) but worsens rollout ranking (Table 2b), likely due to overfitting—where the model learns to predict increasing rewards regardless of input. This overfitting results in 21% inferior relative policy success rate, as seen in Table 3. To avoid overfitting, the main ReWiND model uses only first-frame positional embeddings (Section 3.1.2).
Less Training Tasks. Finally, we additionally check the performance of ReWiND with less training tasks: 5 and 10 instead of the original 20. We can see in Table 3 that the reward model still performs well even with just 5 tasks of training data instead of 20, but policy success rates understandably drop with fewer tasks.
4.4 Additional Analysis and Results
Contact Rich BRIDGE Example.
We additionally plot an example of a ReWiND’s predictions for a contact-rich unfolding towel task from the BRIDGE-v2 dataset [67] contained in the OXE [16] training data used to train ReWiNDin Figure 7. Even though the ReWiND reward function was not trained for many training steps overall in order to prevent overfitting (2k in MetaWorld and 10K in the Real-World, see Section A.1.2), it can perform well on this BRIDGE example. Notice specifically that ReWiND rewards stagnate when the gripper first misses the cloth grasp, and then later increase again as it succeeds on the second retry.
Diffusion Steering.
We conducted preliminary experiments with ReWiND on more expressive policy classes, namely diffusion policies (DPs) [68]. We use Diffusion Steering RL (DSRL) [69] for online RL of DPs. Rather than optimizing the weights of the DP, DSRL runs RL over the latent-noise space of a diffusion policy. This latent space encodes an action prior that RL can refine to improve DPs. The RL policy, conditioned on image and instruction, was trained offline with ReWiND rewards and finetuned online in MetaWorld (Appendix B.2). For seen tasks, this approach improved base DPs: offline RL learned useful priors that accelerated online RL. However, when we attempted to learn new tasks, we found that the action priors learned from offline RL was not representative of the new task as the instruction (OOD and data-sparse regime). We thus trained the latent policy only on images (while keeping the DP instruction-conditioned) As a preliminary result, this DSRL+ReWiND implemented improved DP’s performance on an unseen task 0% 20%. As such, ReWiND is agnostic to policy architectures: single-action-output MLP architectures in MetaWorld, action-chunked transformers in the real world, and now also DPs via diffusion steering.
@@ 5 Limitations
However, the best way to fine-tune these models with online rewards remains an open challenge [72, 73]. One bottleneck is simply that, even with low-rank adaptation techniques that prior work found to help train large policy architectures more efficiently [74, 73, 70], fine-tuning these models takes a lot of compute and real-world time that makes real-world online learning with reward difficult, albeit feasible in simulation [75]. Recent work has also explored how to use RL to improve more expressive policy classes [69, 76, 77]. We show preliminary experiments with Diffusion Steering RL [69] with ReWiND, along with discussions on potential drawbacks to such methods for learning new tasks. We plan to investigate blending ReWiND approaches with such policy architectures and efficient fine-tuning techniques, such as large model steering approaches [69], in the future.
Reward Objective.
One limitation of using video progress as a way to label training data, whether one uses relative targets [13, 41] or fixed targets such as with ReWiND in Equation 1, is that we cannot explicitly incorporate collected failure trajectories into Equation 1 without labels indicating what the final task progress for the trajectory should be. Alakuijala et al. [13] incorporates failure trajectories by using them as contrastive negatives while disabling their sequential ranking loss which relies on frame progress. Similarly, ReWiND can naïvely be trained with collected failure trajectories by only using them as negatives with a progress target of 0 in Equation 1. However, future work can incorporate other objectives, use accurate VLMs, or assume human labelers to label these videos with approximate progress targets.
@@
Another limitation comes from requiring success detection for the reward bonus and terminating policy rollouts upon success. We add a success bonus (detailed in Section 3.2) to account for potential noisy rewards and imperfect success detection by the reward model, given that a human is already monitoring to reset the environment, and we subsequently terminate the rollout upon success. Methods such as those introduced by Ye et al. [47], Zhou et al. [80], Yang et al. [42], which utilize VLMs as success detectors, can remove the need for human supervision during the online phase of ReWiND when combined with reset-free RL. In future work, we plan to investigate ReWiND with reset-free approaches, guided by task performance estimates [81] and automatic success detection for truly autonomous learning.
@@ Acknowledgments
We thank Sid Kaushik for feedback on the final draft of the paper. This work was supported in part by a grant from the Army Research Lab (ARL) Army AI Innovations Institute (A2I2), award number W911NF-23-2-0010. The claims and findings of this work do not necessarily represent the views of the ARL. This work was also supported in part by a grant from the DARPA Friction and Accountability in Conversational Transactions (FACT) program. Additionally, this work was supported by an Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (No.RS2019-II190075, Artificial Intelligence Graduate School Program, KAIST) and National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (NRF-2021H1D3A2A03103683, Brain Pool Research Program; RS-2024-00414822).
@@ A.1.1 Open-X Dataset
We select a subset of datasets from the Open-X Dataset [16]. The subset includes Bridge-V2 [67], BC-Z [82], Fractal [1], CLVR Jaco Play [83], Berkeley Autolab UR5 [84], Berkeley Fanuc Manipulation [85], CMU Stretch [86, 87], Stanford Hydra [88], UCSD Kitchen [89], Austin BUDS [90], Austin Sirius [91], and DROID [92]. These datasets were selected for their high-quality, task-oriented manipulation trajectories (i.e., no play data or extremely high-level annotations). These datasets provide around 350k trajectories and 58k total unique task annotations. To ensure meaningful trajectories for training the ReWiND reward model, we postprocess the data to remove trajectories with less than 5 timesteps. We subsample the videos in the datasets to 16 frames for reward model training, as we did not see a noticeable benefit from training it with longer videos.
@@ A.1.2 Reward Function
The cross-modal sequential aggregator takes these tokens as input and produces a per-image embedding used by an MLP to produce per-timestep rewards. The cross-modal sequential aggregator is a causally masked transformer (PyTorch nn.TransformerEncoder) composed of 4 layers, each with 8 heads with a combined hidden dimension of 2048. We add a learnable positional embedding to only the first frame of the video sequence embedding. In the ReWind reward function training phase, we trained 2k steps for Meta-World and 10k steps for Real-World robot experiments, with a batch size of 1024. Each batch includes 80% data from and 20% target environment data from . Each video in the batch has an 80% probability of having video rewind augmentation, and independently, a 20% percent probability of having a mismatched video-language pairing with 0 progress target (see Section 3.1). For the rewind augmentation, the split point is uniformly sampled from the video clip, and the rewound sequence is obtained by taking all frames from the split point to the end of the clip. In order to better simulate policy execution videos that look almost successful, 10% of the rewound videos will only have their last 3 frames rewound. No extensive tuning was performed on these per-sample rewind and mismatch probabilities; they were heuristically chosen during initial small-scale experimentation and then fixed for all experiments.